OpenStudy (anonymous):

write an equation that shows when 4 is taken from 3 times a number the result is 14 solve

OpenStudy (anonymous):

3x-4=14

OpenStudy (anonymous):

3x - 4 = 14 x = 6

OpenStudy (anonymous):

3x-4=14 to solve: add 4 to both sides: 3x=18, then divide both sides by 3 and get:x=6

OpenStudy (anonymous):

shouldn,t it be 3 X a=14

OpenStudy (anonymous):

it's an algebra question

OpenStudy (anonymous):

well since algebra starts with the letter "a", we can use "a" as a variable, but the way we express the problem in numbers is the same: 3a-4=14

OpenStudy (anonymous):

oh the answer stay the same thank

OpenStudy (anonymous):

yes a=6